Weekly report of problems
swings at za=10 with the tilt sensors were done to try and measure
any residual platform tilts. The Pointing checks showed a 9 asec 1az
that could come from a platform tilt. The tilt sensors did not measure
any residual tilt. The 9 Asecond tilt is probably true but that would
a .0025 degree 1az term. The tilt sensors fits usually give rms fits
UT1 - UTC dec06 thru Jun07.
11dec06: fit 1 az and sin(za) terms to pointing errors and update
the pointing model.
phone detector birdie at 1548 Mhz Data taken back on 08nov06
that the new cell phone detector has a birdie at 1548 Mhz that is about
15 times Tsys in a 390 hz channel width when the cell phone detector is
in the parking lot adjacent to the control room.
26oct06: 430 transmitter had rf on before ack light active. Problem was
the monoplexor in the ch was taking 3 useconds longer than the
in the carriage house. The delay was initially measured and the
limits set to 12 usecs from tx ipp to rf on. When the software limits
set i think it took 3 usec (signal up) 3 usecs turnon, 3 usecs (signal
down) = 9 usecs. This gave 3 usecs of slop for the software
With the added 3 usec delay of the ch monoplexor, it is now possible
the rf request to come before the ack. Jon's box will hold off the rf
the ack gets back.
az1 encoder failed from lightning. Problem was the encoder board
cable test upstairs iflo showed loose cables in 750 mixer chassis.
28sep06: added alfa
stability page. Uses TsysPolA/TsysPolB for a2010 drift scan data.
months aug06, sep06 (thru 20sep06). Shows when and how much beam 5b and
stability page. Added months jan06 thru jun06.
03oct06: The alfa
stability monitoring (polA/polB) for sept shows that different
were drifting and jumping during the month:
beam 5B was jumping around on (03,09 thr 14,) sep06. It was especially
bad on 9sep and 12sep. On 11sep the jumping started. The 12 second
birdies/compression seemed to be magnified on 11sep06.
6B was moving on 05sep.
3A,4A,5A drifted together on 16-17 sept and30 sep. This is also seen by
looking at the first
ten strips of each day for sept for a2010 on these days.
09aug06: power dip caused dewar monitoring to hang up. Had to cycle the
power on the gpib to ethernet converter up in the rfi box in the
The ups for this box is probably not working very well
stability of cbh using the wapps and 100 Mhz bandwidth, Compare
with the data taken with square law detectors.
10aug06: dewar monitoring stopped working after a power dip. Brought
down to control room to test it out. On 11aug06 it started working
with a 4 % packet loss. It was using fiber c1-11 and c1-12 (main
cables rcvr room to dome tx). We found that main cable c1-11 had 5 db
loss than c1-12. We switched c1=11 to c1-13 (which was unused. It
used to be the dome velocity monitor of hagens.). Still had 4% packet
(but we left it connected to c1-13 ...c1-11 is now unused). Dewar
system is running in normal config with downstairs repeater installed.
The 4% packet loss is more than normal (should be 0). It may be a xcvr
or xcvr power supply problem upstairs or downstairs. Should check the
connection in vertex shelter with pc to see if its packet loss is
12aug06: td12 encoder stopped working around noon after a power dip.
encoder always read 9" (without giving an error). When the td was asked
to move, it would move up or down till it hit the limit switch (since
was looking at the encoder for the position feedback). I manually
td12 to 12" (using a tape measure) and arranged the other 2
to be at the right position. I checked the voltage on the encoder board
and it was ok (5 volts). The tiedowns should be left off till this is
430 yagi performance after moving the antenna closer to the line feed
(before the low loss cable was installed).
07jul06: Tsysall would not work. Upstairs power meter has been shut
Someone had been working in the rfi shielded box where the power meter
was installed. When the reinstalled the box cover, it pushed the off
of the power meter. After finding the problem we moved the power meter
farther into the box to not colide with the cover.
near 1320 Mhz appears in a2049 data on 06jul06.
focus curves sbtx old and new horns.
16may06: for IF monitor display added the alfa filter bank. It will say
fbAlfa wideband or fbAlfa 1390-1490.
1gb ethernet switch causes rfi that can be seen in the 327 receiver.
30apr06: power failure caused little star in the turret box to lose its
program. Had to reload the program.
at the new ac units to see if the 1 Mhz comb is still present.
the old Ac unit on and off to make sure we can see it.
15feb06: joanna rankins had problems with her polarization data
on 26nov05. When i look at her data, i see no correlation in the
polA, polB cal values (as if it was taken with 2 diodes rather than 1).
We were running with the single pol B diode till 23nov05. So the new
were installed when she took her data. It may be that the software that
was forcing the polB 90deg phase shift was not updated till 28nov05
If she took her data with the 90deg phase shift it still seems like
should be correlated power (but maybe not in the cross spectra
i was looking). Need to understand this a little better.
21jan06: It is no longer possible to send the carriage house signals to
the wapps. This caused a problem with dan stinebring when he tried to
50 Mhz, 430 Mhz data simultaneously. The loss of functionality occurred
when bill sisk removed the external inputs of the buffer amplifiers.
23jan06: The 50
Mhz receiver has a frequency offset of about 150 hz when tested
the birdie maker. The synth used to generate the frequency was checked
and it is locked to the station clock.
02dec05. galfa observers had reported at ripple in beam 2a for the
band spectra. The wide band spectra did not have the ripple. On 02nov05
we switched the galfa 2a,2b inputs. The ripple stayed in pixel 2a so
problem was in the digital card (we then returned the cables to
normal order). The system was rebooted by cycling the power (on the
supply). After this the problem went away.
05dec05: A2125 was doing mapping with the 430 Mhz reciever. Periodic
with periods of 1 minute were evident in their data. On 02dec05 i took
2 hours of data to characterize the
27oct05 lbw filter bank. PolB 1230 to 1470 Mhz filter (filter
3) is drifting in gain. The polA filter is ok. Using some of the other
filters in the filter bank shows no drift. So the problem must be
the box with one of the connections for that filter.
29oct05: murray sees calibration
jumps in his OH data aug-oct05.
430 carriage house cal values. Looks like there is a bad match in
starting at the diode. The reflection is 3.75 meters in air, (2.75 m if
velocity is .7c).
30sep05: at midnite (start of 30sep05) a 1 second tick was
inserted into the station clock. The clock was off by 1 second (The
clock provided a time that was 1 second behind the correct time).
offset was corrected about 8:30 am 30sep05. During the morning of
30sep05 the tiedowns were acting strange and the rotatary floor
a bit. I think this was because of the 1 second time difference. The
star computers did not jump with the 1 second tick insertion. The time
stamped positions that they received were one second late, so most of
tracking commands sent to them would have be ignored.
03oct05: Early morning, the sband radar run had trouble with the cal
that fires the cal. This program runs on the corp1 computer. Programs
not able to communicate with it. Rebooting the cor crate fixed it. The
backplane network for the correlator crate has been subnetted. This
has occurred in the past. That time a computer using the "other" part
the subnet sent arp packets saying that they also belonged to the "cor
backplane" part of the sub net. This caused requests to go to this
computer rather than the cor backplane (there is no evidence that this
is what actually happenned this time..).
07sep05: dewar monitoring failed 06sep05 after power dip. Had to cycle
the power on the upstairs fiber to gpib transceiver. The fiber repeater
in the racks downstairs for the platform ethernet is not on a ups. This
is probably causing the problem. Luis put the repeater on the ups in
07sep05: tension in tiedown cables after td4, td8 were recalibrated on
02sep05. Td4-2 looks like it's gain needs to be recalibrated.
12sep05: 1 section of rolling surface of turret floor replaced
On 12sep05 did two
turret floor spins with the tilt sensor to see if there were any
when the old to new rolling surface passed over a roller.
31aug05: 34970 multimeter for dewar monitoring continued to fail.
the device down to the labs. Had trouble turning it on with no cards
in. Will try the spare 34970 upstairs tomorrow.
31aug05: looks like lbw
tsys has jumped up 4 or 5 K after the 26aug05 warmup then
01sep05: replaced the 34970 with the spare and the dewar monitoring now
works (used the old plugin cards).
02sep05: replace spectrum analyzer in rfi shack with spectrum analyzer
that was in the ant test range. rfi monitoring working again.
02sep05: td 8 and td 4 the amplifiers for 4 loadcells had their
regulators broken. replaced the voltage regulator chipes and
the loadcells. tensions on tiedowns now working. The tension on td 4
2 is drifting by a few kips (with no load present). The db25 filter for
td8 was bad. It was removed (we need to order more of these). td is
missing the db25 fitler.
02sep05: lbw is only using the 2nd diode (since diode 1 was replaced
the spare diode has not been calibrated). This configuration will
not work with the lbw hybrid in to give circular polarization. When
2 feeds both pols, there is a 90 deg phase shift between polA and polB.
When this is passed through the hybrid to create the circular, no cal
gets to one of the circulars. Use linear pols until we fix diode 1.
02sep05: p2030 alfa pulsar search. Observers complained about large
in data. Showed them that it was from the out of band faa radar.
04sep05: td12 db25 filter for the pcu shorted out the 24 volt supply.
done with tiedown tilted by about 8 inches (need to check the pointing
error versus tilt angle).
05sep05: 16khz comb at 327Mhz. Grounded
the motor shields 29aug05. Improved things for awhile but then it
06sep05: lbw increase in Tsys is coming from the
resonance at 1400 Mhz. The 27aug05 problem of murray lewis with no
signal in polB was caused by the h90cal and his selection of the
polarization option. The signal was present, but no cal got through to
one polarization (because of the h90cal 90degree phase shift and the 90
degree phase shift of the hyrbrid to create circular).
tsys for mathews feed (22aug05). Tsys is about 151 K. There
a 220 second ripple in the cal
27aug05: lbw . murray lewis took some oh data in the morning and
he got no signal in polB. He was using the hybrid and the 1550-1820
The upstairs power meter showed power in both pols (-40 dbm). The
got no data in polB. Later users of lbw had not trouble (although they
had slightly different setups). The downstairs synths are not (rf off)
since they are shared between the two pols.
27aug05: the 16
Khz comb in the 327 Mhz receiver is coming from the alfa
29aug05: td8 . power cycled on td8 around 3pm. Later that evening (8pm)
it was noticed that the encoder readout value for td8 was not changing
(even though the jack was moving). Cycled the power on td 8 (ups and
and the encoder started reading correctly.
30aug05: on 26aug05 about 14:00 the rfi hilltop monitor stopped
Turns out the spectrum analyzer is no longer communicating on the gpib
bus. Brought the analyzer down to the lab.
30aug05: td8 has been reading motor over temp. The circuit for this bit
was wrong. Updated the circuit to read correctly. The other tiedowns
the same circuit. they were updated to the new version.
30aug05: the power cable for the 34970 used for the dewar monitoring
bad. Moving the cable caused the power to the device to go off.
the cables but it failed about 15 minutes later.
18aug05: lightning smoked td 12. Tensions went bad for awhile and then
the encoder position went to 24Inches. This usually means the little
is not reading the encoder values correctly. When jorge,luis looked at
the little start, the front panel display was blank. Could not change
star since lightning started up again. will do it tomorrow.
19aug05: td 12 started working after encoder interface board was
Looks like the kips on once load cell of td 8 are now off by 20 kips.
occurred during the lightning ...
20aug05: datataking hung up after the power changeover from generator
comercial saturday evening. All 4 of the downstairs synthesizers had
their power. You could see this since all 4 of them had the rf off
set. After the power cycling the gpib bus that controls them was hung
I cycled the power on the 4 synths and then restarted the if2Prog that
controls them. These synthesizers need to be put on a ups (if
are currently on a ups, then the battery/load needs to be checked). The
observer probably lost over an hour of observing time...
21aug05: the galfa piggy back on alfalfa (they're calling it togs)
of an rfi at 1421.6 Mhz. This is an intermod of the 1350 faa radar
in the galfa mixers. 1350 is at 285If1, lo1 is 195.845. The intermod is
3*192.845 - 2*285= 17.53Mhz at baseband. baseband dc is 1439.16 and
flipped so this comes out to (1439.16-17.53=1421.63). Alfalfa is
not going to move their sky center frequency. I guess if we have a
bank before the galfa mixer we could cut the 1350 radar...
10aug05: problem with distomat 1 was the fiber connection. Fixed it and
the distomat started working again.
12apr05: during 430 run, saw interference for awhile. Some of it was a
2Mhz comb. Was not able to trouble shoot since radar run was ongoing.
away after awhile.
03aug05: distomat 6 has been failing more often then the other
(see distomat failures for jul). Electronics checked all 6 distomats
cleaned the windows (when distomat 6 was cleaned, the coating on the
of the window slid off when glass cleaner was a>
ing was for??). After this, distomat 6 performed the same as the other
- 04aug05: Observer using cband high had intermittant ripples in
power. Period of 5 hz. Lasted for a few minutes and then stopped.
2 days in a row. Amplitude was a few % of Tsys. When looking at the
saw a ripple in polB of about 1% of Tsys. Checked the cables and found
2 bad cables: dewar to postamps polA, and postamps to If/Lo cabinet
Replaced both cables. Vibrations in the system were probably causing
03-04aug05: maintenance is installing some beams in the feed tower to
some beams that are deflecting because of the kevlar cables. This may
the pointing ??? need to check. Took some calibration data 05aug05
and 08aug05 cbhi. The pointing
did not look that bad. There may be a small ramp in azerr,zaerr vs
za but it will take more data to nail it down.
05aug05: td12 trouble. motor disabled with overspeed condition. Mainly
occurred when trying to pull down. Jorge went down and turned the cmd
pot 1/8 of a turn and it started working. There is something funny
on with this td. Also found that the jack was full of water.. would
it out when you pulled the jack down to around 18 inches. Maintenance
they would look into trying to seal it better.
05aug05: pulsar observers complained that alfa was a lot less sensitive
than lwide when looking at known pulsars (they were using pixel 0).
Tsys for alfa and it looked ok (see alfa
tsys by month). They did say that the 12 second radar may have been
affecting their ability measure the sensitivity.
05aug05: tim hankins complained that he didn't see the crab pulsar
Also saw a large 1.6 second ripple in the total power (probably a
few % of Tsys). Saw it in cbh and in sbh. Could be pointing error since
we are looking at the crab nebula?? took some pointing data. Took some
data on 05aug05 (sbhi) to check the pointing and on 08aug05 (cbhi) to
the system performance. The measurements
showed that cbhand high was working ok.
09aug05: tim hankins saw crab pulses at cbhi and xband.
09aug05: 8pm. distomat 1 failed. remained off till next day.
29jul05: dewar monitoring failed again after a power dip. They went up
and checked and the ups has a problem. This needs to be checked out
it also holds up the monoplexor driver for the 430.
29jul05: still problem with td12. overspeed.
28jul05: td 12 failed after power dip. Changed little star, dac, signal
conditioning. Still have problem with overspeed. The signal
board might not be configured correctly. There is confusion between the
differential version (old ones) and the single ended version (new
Not obvious what the old configuration was.
28jul05: dewar monitoring stopped after a power dip. Jorge and luis
up and found that the power strip in the rightmost rack in the rotary
room is not on a ups. After talking to tony noya he told me that only
leftmost rack is on a ups. The rfi box in the right most rack contains
the power meter used to measure the if1mp power for the fiber optics.
also holds the dewar montoring. The power dip was screwing up the
transceiver for the dewar monitoring. Luis connected the power strip of
the rightmost rack to the ups that is used for the monoplexor driver.
explains why we have always had problems with the dewar monitoring
a power dip. I had always thought that both racks were on a ups. The
rack also contains some fiber transcievers and the 20 Mhz buffers. I
know whether they were also plugged into the power strip in the
rack, or they had power cables that went to the left rack. We need to
that the 2nd ups that we plugged into can hold the combined
strip load. We checked and it holds the power strip load just find (but
the monoplexor was off).
29jul05: power dip and dewar monitoring hung up again. Since rfi box
on ups, problem must be in the fiber ethernet repeater in the
room upstairs. The repeater was put on ups 01aug05.
01aug05: People reported that pixel 2b in galfa has it's power moving
again. Drifts off for awhile and then comes back. This sounds like the
with pixel 2b.
01aug05: p2030 alfa pulsar search saw large 12 second compression in
data. Turns out to be compression
from the faa radar.
02aug05: Upstairs if/lo, user selected the straight thru option (if1
5"). This should bring the rf down the fiber (no 1st lo mixing). The
appeared on polA, but their was no signal on polB. Checking the
of sw3b,sw5b upstairs, the lights had position 5 selected correctly.
out that someone had taken sw3b, pos5 and routed it to an unused
The signal was not passing thru the amplifier. This must have happened
when someone was doing a test upstairs and didn't reconnect the cables
correctly. Plugged the cables in correctly and there is now a signal in
both polA, polB for the straight thru (ifnum 5) position.
13jul05: The power
level in pixel 2b of the galfa spectrometer is drifting/jumping.
taken at the same time with the wapps shows no drift/jumps. The cables
for galfa pixel 2b were all tight.
14jul05: The dewar monitoring gpib-enet converter was replaced. The
that broke was a 10Mbit gpib-enet. We replaced it with a 100Mbit
I had tried to use the 100Mbit device before with vxWorks and
never get it to work. It turned out that the programs i had to
the ipadress on the 10Mbit device did not work for the 100Mbit device .
Jorge downloaded some windows software and got the addresses setup ok.
This means that we can use the 100Mbit devices with the vxWorks
(if we want to).
14jul05: a2055 lost about half there run. Gui said waiting 86402
to be on source. The basketweave routine had trouble crossing midnite
the pattern started too close to midnite. Wrote some tcl code to fix
problem. Arun was going to install it when he got a chance...
15jul05: During the day the tiedowns were disabled because people were
working with the distomat temperature sensor. The log book says
the tiedowns should be turned off till further notice. I think
was a mixup in communications. After working on the distomats there was
probably a thunder storm and the person probably told the operator to
the tiedowns off "until the weather clears up.." .. Anyway the tiedowns
were left off for the entire weekend. On monday morning i turned
them back on.
pixels 4a and 6a were interchanged. This occured at the output of
distribution amps in the control room (the wapp data was not affected).
The pixels were swapped on 11jul05 and were put back in the correct
06jul05: cima was configuring the cal multiplexor to fire the 25Hz cal
(if2_calsrc 2 ) rather than the interrim correlator (if2_calsrc
When the interim correlator was initialized by cima.
07jul05: The alfa
Pixel 4A 174 Khz ripple is coming from the fiber receiver.
replaced the pixel 4a fiber receiver for alfa and the 174 Khz
went away. Pixel 2a still has some ripple.
09jul05: dewar monitoring stopped working. The upstairs fiber receiver
did not have any lights on. The problem is either the fiber receiver or
the gpib to ethernet converter.
09jul05: the distomat temperature sensor has been failing
This has been going on for awhile.
12jul05: lbw channel B gain or tsys was changing with time. The
and 1320 HiPass were in. When the 1370 hiPass was taken out, it
more stable. 13jul05 found that the digital line that controls the
for the 1370 hipass polB was broken. Fixed the cable.
dewar warms up during sband radar run.
01jul05: Looked at the 174 khz ripple in pixel 4a of alfa. (more
05jun05: Td12 is not starting up consistantly. The commanded power on
uses a two step process (see td
power on sequence). The first power1 sent by the little star is not
always sending the 24 volts to close the master contactor relay and
power to the pwrsup/amp. Tracing the cables it looks like there is a
cable and the signal drive is floating..
22jun05 - cband polarization switch hooked up. cbpol (lin,cir) switches
it back and forth.
23jun05 - took on/off cals of lbw at 1400 Mhz to see the resonance in
25jun05 - tower 12 hit by lightning. Td 12 releases, stays released for
about 1 day (more
17jun05 - The old method of computing the cals would find a single
for each 25 Mhz band and then fit an Nth order polynomial across the
band (1120=1720 for lbw). This polynomial was then evaluated at 25 Mhz
steps and placed in a lookup table for later use. The lbw cal data had
lots of rfi (on the sky) and a large ripple in polA with a period of
75Mhz. The old method did not do a very good job of fitting the data so
i switched to a new method that fit the entire bandpass at once.
18jun05 - distomat 4 stopped working at 13:00 (sat). Stayed off the air
until 20jun05 (10am monday). Luis murray went down and cycled the power
on the main box for distomat 4 and it started working. There was no
storm at the time it failed since the other distomats remained
If all that is needed is to cycle the power on the distomat, then this
is something that an operator could do. It probably would be a good
for someone to understand why cycling the power should make a
For 2 days the tracking of the telescope was not "optimal" since the
switched to using the temperature sensor (and model) rather than the
21jun05 - transfered power from generator to public power in the
After the switch, the 4 synthesizers that drive the iflo (the original
4) would not function. The "rf off" label was selected which implies
the synthesizers saw a power dip. Turning the rf back on made no
The gpib controller could not communicate with the 4 synthesizers.
all 4 of the synthisers off and then on caused them to start working
It looks like the ups for these 4 synthesizers is not holding up under
the power dip. The synths are getting a power glitch and they are
back in a funny mode where the gpib communications is hung up. We need
to check the ups to make sure it is large enough (and that it's
21jun05 - finished reducing the lbw cals. Installed them in the
area and the idl data directory. The cal for polA looks like it
more ripple than it used to (about a 1.2 meter cable).
21jun05 - polA lbw has a resonance at 1400 Mhz. It is about 30% larger
than tsys. The tsys values that have been taken are reading 35 K.
to 1403 Mhz, the Tsys is only 29-30K. The
10jun05 - measured lbw cals sky/absorber. Saw a punta salinas
radar around 1305 Mhz. Punta salinas says it was not them and it was a
bit off of their frequency. Also took cband absorber data for cband
(so cheo didn't have to come in monday morning at 6am).
11jun05 - riccardo reported large birdie (alfa) at 1428Mhz for the
a2010 data. (see 1428
Mhz birdie). I've asked riccardo to send me the scan numbers for
bad data. When i look at it I see the large birdies but i don't
it spreading over all of the band.
11jun05 Riccardo reported a wandering birdie n alfa in jun02 data. I
and it is there. Need to investigate this when we get some telescope
13jun05 - took cal lband sky data over again since radar at 1305
was gone. Also took lband data to measure ratio of other cals to High
cal. Took cband sky data with high correlated cal (but it was cloudy
drizzly so probably not that great). Had no time to do the ratio of
high correlated cal to the other cals. Looked at lbw cal data and the
data was a bit unstable at the beginning. Probably because the absorber
was removed from the air conditioned room right before the measurement
started. The cal measurements need more that 2 hours to get the most
of the time (so we can to cal, absorber and measure the ratio of
to other cals at the same time ... and so phil doesn't have to get up
5:30 am so often).
13jun05 There is something funny going on with the lbw cal data. The
cal value for pol B increased by about 1.5 Kelvins after the polB cal
was fixed. This makes sense and means that the cable was really broken.
The Tsys for polB using the old cal values and the new cal deflection
35 K where it used to by 28 K. If the amount of power in the cal
has increased and i am using a smaller value for the cal in K then the
Tsys should be less. I did notice that there seems to be a bump
polB at 1400 Mhz that is about 30% of Tsys and about 2 Mhz wide. This
to be looked into.
13jun05 - installed cband pol switch software. tried testing it but it
turns out the cables haven't been connected.
occurence of rfi for may05 updataed
03jun05: 1am nevis server fails. This is the yp server (providing info
on discs, users, permissions,etc for the entire network). There is a
server but the computers didn't switch to it. Datataking
hangs, telescope offline.....Spent 1am thru 8am with arun (on the phone
from washington dc) getting services moved around so things work.
observer got back online about 7am. Arun then worked on the mailserver
while i went to bed. Manny found that the problem with nevis was
that the 2 cooling fans in the computer had stopped so looks like cpu
05jun05: updated a2010
monitor data for may05. The problem with the faa radar causing
jumps in the total power and then a long recovery time (many seconds)
still a problem. This is also causing a problem with the alfa continuum
group (desh, chris, russ taylor).
comb in 327 receiver. problem with P1979,a2057. Still looking into
26may05: alfa pixel4a still has 173Khz ripple with amp of
28may05: lbw stokes calibration table included data for 4
taken jul04,sep04,oct04,apr05 (more
info). The delta G parameter shows that the cal values for lbw are
not constant. The drift is probably in polB. The yearly plots of
TsysA-tsysB also shows a drift in Tsys (TsysA-TsysB
(see lbw 2005)).
30may05: distomat 4 loses power 18:00. next day willie finds
it is an ac power filter at the input of the distomat canister that is
blown (it wasn't the surge protector after all). He didn't have
to change it because of the sband transmitter.
30may05: chris salter reports a 16 khz comb at 327 Mhz. I
data on 31may05 and saw the comb close to the 15.85 khz that was
from the cameras 13oct03 (15.85
Khz comb from cameras). We couldn't easily test the cameras because
the switch to power off the cameras in the control room has not been
for a long time. They had trouble making the relay upstairs work so it
was bypassed and power was provided directly to each camera. We need to
get this fixed. The operators new about this, but it was never
to me, rey velez, or anyone one else in electronics (that i am aware
31may05: galfa pixel6a and pixel4a cables were switched at
of the receiver room distribution amplifiers. It has been switched
at least 02may05. It was not switched on 12dec05 (i've got no data
these two dates). It was fixed at 11am 31may05 (more